# A Hybrid Systems Model of Feedback Optimization for Linear Systems: Convergence and Robustness

### v4

@@ V-C Complete Hybrid Convergence
Then using (59) in (58) and combining sums of integrals into one integral as was done below (53), we have
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\int_{0}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|q^{\frac{\ell-1}{2}}d_{\mathcal{U}}\\ \cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\int_{0}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}.$ (60)
Next, we observe that
@@
where the inequality follows from Theorem 2 in [28, Chapter 1.9].
@@
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|q^{\frac{\ell-1}{2}}d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\cdot\\ \|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|.$ (63)
@@
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|q^{\frac{\ell}{2}}d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big].$ (67)
@@
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)(\|\phi(0,0)\|_{\mathcal{A}}+r)\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}[1-\exp({-\rho}t)]\\ +\frac{M\left\|B\right\|q^{\frac{\ell}{2}}d_{\mathcal{U}}}{\rho}[1-\exp({-\rho}t)]-r,$ (70)
@@
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)\|\phi(0,0)\|_{\mathcal{A}}\\ +(M-1)\left(\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}+\frac{M\left\|B\right\|q^{\frac{\ell}{2}}d_{\mathcal{U}}}{\rho}\right)\exp({-\rho}t).$ (71)
@@
In the limit, the state $x$ is asymptotically no farther than a distance $r$ from $\tilde{x}$. The quantity $r$ can be reduced with a pre-feedback controller that is applied before feedback optimization is used. More precisely, if the pair $(A,B)$ is controllable, then using pole placement we can use any $\eta>0$ to select $\rho$ such that $\rho\geq\frac{M\left\|B\right\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{\mathcal{U}}}}{\eta}(1+q^{\frac{\ell}{2}})$. This $\rho$ ensures $r\leq\eta$ and hence that $\|x(t,j)\|_{B_{\eta}(\tilde{x})}\to 0$ for any desired $\eta>0$.
@@ V-D Global Convergence and Robustness
Consider the hybrid system $\mathcal{H}_{FO}$ from (17) and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (8), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$, where $Q_{u}$ is from (8) and $L:=\lambda_{max}(Q_{u}+H^{\top}Q_{y}H)$. For each maximal solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)$, for each $(t,j)\in\textnormal{dom }\phi$,
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M^{2}\exp(-\rho\tau_{c,\min})\exp(-\rho t)\left\|\phi(0,0)\right\|_{\mathcal{A}}\\ +\frac{M^{2}\left\|B\right\|d_{\mathcal{U}}}{\rho}\exp(-\rho t)\times\\ \Bigg[\left(M\exp(-\rho\tau_{c,\min})-\frac{1}{M}\right)\Big(1+q^{\frac{\ell}{2}}\Big)+2\Bigg],$ (73)
@@
which happens precisely because no computations have been performed to change the input. This case is captured by using the same steps to reach (67) but with $\ell=0$, which gives
$\left\|x(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq\\ M\exp({-\rho}t_{\bar{\alpha}(2)+2})\left\|x({0},{0})-\tilde{x}\right\|+\frac{2M\left\|B\right\|d_{\mathcal{U}}}{\rho}.$ (75)
To maximize the upper bound, we observe that
$\tau_{c,\min}\leq t_{\bar{\alpha}(2)+2}$ (76)
since, by the point at which flow time $t_{\bar{\alpha}(2)+2}$ is reached, the timer $\tau_{c}$ has undergone at least one reset to the interval $\tau_{c,\min},\tau_{c,\max}]$, followed by $\tau_{c}$ reaching zero. Using (76) in (75) gives
$\left\|x(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq\\ M\exp({-\rho}\tau_{c,\min})\left\|x({0},{0})-\tilde{x}\right\|+\frac{2M\left\|B\right\|d_{\mathcal{U}}}{\rho}.$ (77)
Then, using (68) and (69) with $(t,j)=(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)$ and combining like terms gives
$\left\|\phi(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)\right\|_{\mathcal{A}}\leq M\exp({-\rho}\tau_{c,\min})\left\|\phi({0},{0})\right\|_{\mathcal{A}}\\ +M\exp({-\rho}\tau_{c,\min})\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(1+q^{\frac{\ell}{2}}\Big)\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(1-q^{\frac{\ell}{2}}\Big).$ (78)
After hybrid time $(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)$, there are at least $\ell$ gradient descent iterations performed before each subsequent change in the input. Then Theorem 1 applies but with the upper bound on $\|\phi(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)\|_{\mathcal{A}}$ in place of $\|\phi(0,0)\|_{\mathcal{A}}$. The result follows from making that substitution.
@@
$C_{\rho}:=\{\zeta\in\mathcal{X}:\tau_{c}\in[0,\tau_{c,\max}+\theta_{c,\max}],\\ \tau_{g}\in[0,\tau_{g,comp}+\theta_{g,comp}]\},$ (79)
where $\theta_{c,\max}\in(-\tau_{c,\max},\infty)$ and $\theta_{g,comp}\in(-\tau_{g,comp},\infty)$. These perturbations allow $\tau_{c}$ to take values larger than $\tau_{c,\max}$ and similar for $\tau_{g}$. Jumps are still triggered when at least one timer reaches zero and hence we use $D_{\rho}:=D$ in the perturbed system model.
@@
$F_{\rho}(\zeta):=\left(\begin{array}[]{cc}\big(A+\hat{A}\big)x+\big(B+\hat{B}\big)u\\ 0\\ 0\\ 0\\ -1+\kappa_{c}\\ -1+\kappa_{g}\end{array}\right),$ (80)
@@
For $G_{1,\rho}$ we have
$G_{1,\rho}(\zeta):=\left(\begin{array}[]{cc}x\\ u\\ y_{s}\\ \Pi_{\mathcal{U}}\left[z-\gamma\nabla_{u}\Phi(z,y_{s})\right]\\ \tau_{c}\\ \tau_{g,comp}+\theta_{g,comp}\end{array}\right),$ (82)
where $\theta_{g,comp}$ is from (79). This perturbed jump map allows for $\tau_{g}$ to be reset to values other than $\tau_{g,comp}$ after a gradient descent iteration is performed. For $G_{2,\rho}$ we have
$G_{2,\rho}(\zeta):=\left(\begin{array}[]{cc}x\\ z\\ \big(H+\hat{H}\big)u+d\\ z\\ \left[\tau_{c,\min}+\theta_{c,\min},\tau_{c,\max}+\theta_{c,\max}\right]\\ \tau_{g}\end{array}\right).$ (83)
Here, $\hat{H}\in\mathbb{R}^{p\times m}$ models perturbations to the matrix $H$, including those that come from the errors $\hat{A}$ and $\hat{B}$ as described above, as well as errors in the output map $\Psi$. The interval to which $\tau_{c}$ is reset is perturbed with constants $\theta_{c,\min}\in(-\tau_{c,\min},\infty)$ and $\theta_{c,\max}\in(-\tau_{c,\max},\infty)$ that satisfy $0<\tau_{c,\min}+\theta_{c,\min}\leq\tau_{c,\max}+\theta_{c,\max}$, which ensures that $\tau_{c}$ is reset to a non-empty set, though both endpoints of the interval can be perturbed. Finally we have $G_{3,\rho}(\zeta):=G_{1,\rho}(\zeta)\cup G_{2,\rho}(\zeta)$. We define
$\rho(\zeta)=\max\{\theta_{g,comp},\|\hat{A}x\|,\|\hat{B}u\|,\|\hat{H}u\|,\\ \kappa_{c},\kappa_{g},\theta_{c,\min},\theta_{c,\max}\}$ (84)
@@
$\mathcal{H}_{FO}^{\rho}=\begin{cases}\dot{\zeta}\in F_{\rho}(\zeta)&\zeta\in C_{\rho}\\ \zeta^{+}\in G_{\rho}(\zeta)&\zeta\in D_{\rho}\end{cases}.$ (85)
@@
Definition 2 ($(\tau,\epsilon)$-closeness [18, Definition 5.23]).
@@
Consider the hybrid system $\mathcal{H}_{FO}^{\rho}$ with $\rho$ as defined in (84), and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (8), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$. Then, for every $\epsilon>0$ and $\tau>0$, there exists $\delta>0$ with the following property: for every solution $\phi_{\delta}$ to $\mathcal{H}_{FO}^{\delta\rho}$, there exists a solution $\phi$ to $\mathcal{H}_{FO}$ such that $\phi_{\delta}$ and $\phi$ are $(\tau,\epsilon)$-close.
@@
Since $\mathcal{H}_{FO}$ is well-posed and its maximal solutions are complete from all initial conditions, the result follows from [26, Proposition 6.34]. ∎
The interpretation of Theorem 3 is as follows. Let a time $\tau$ and an error $\epsilon$ be given. Then there is a nonzero perturbation $\delta\rho$ that we use to get the perturbed system $\mathcal{H}_{FO}^{\rho\delta}$, and any solution $\phi_{\delta}$ to $\mathcal{H}_{FO}^{\rho\delta}$ is $(\tau,\epsilon)$-close to some trajectory that the unperturbed system $\mathcal{H}_{FO}$ could have produced.

### v5

@@ V-C Complete Hybrid Convergence
Then using (54) in (53) and combining sums of integrals into one integral as was done below (48) while separating out the $p=0$ case, we have
$\left\|x(t,j)-\tilde{x}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\int_{0}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1}-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}\\ +\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|q^{\frac{\ell-1}{2}}d_{\mathcal{U}}\\ \cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\int_{0}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}},$ (55)
where the first integral does not have a $q^{\frac{\ell}{2}}$ term to account for $\alpha(-1)$. Next, we observe that
@@
where the inequality follows from Theorem 2 in [23, Chapter 1.9].
@@
$\left\|x(t,j)-\tilde{x}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}q^{\frac{\ell-1}{2}}}{\rho}\big[1-\exp({-\rho}(t-t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1}))\big]\cdot\\ \|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}q^{\frac{\ell-1}{2}}}{\rho}\big[1-\exp({-\rho}t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1})\big]\cdot\\ \|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|.$ (58)
@@
$\left\|x(t,j)-\tilde{x}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}q^{\frac{\ell}{2}}}{\rho}\big[1-\exp({-\rho}(t-t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1}))\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1})\big].$ (62)
Next, using $\tau_{c,\min}\leq t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(1\right)}+1}\leq\tau_{c,\max}$ and $\exp({-\rho}(t-\tau_{c,\min})=\exp({\rho}\tau_{c,\min})\exp({-\rho}t)$ gives the upper bound
$\left\|x(t,j)-\tilde{x}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}q^{\frac{\ell}{2}}}{\rho}\big[1-\exp({\rho}\tau_{c,\min})\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}\tau_{c,\max})\big].$ (63)
@@
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)(\|\phi(0,0)\|_{\mathcal{A}}+r)\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}q^{\frac{\ell}{2}}}{\rho}\big[1-\exp({\rho}\tau_{c,\min})\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}\tau_{c,\max})\big]-r,$ (66)
@@
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)\|\phi(0,0)\|_{\mathcal{A}}\\ +\frac{M^{2}\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(2-\exp(-\rho\tau_{c,\max})+q^{\frac{\ell}{2}}\Big)\exp({-\rho}t)\\ -\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(1+q^{\frac{\ell}{2}}\exp({\rho}\tau_{c,\min})\Big)\exp({-\rho}t)$ (67)
@@
In the limit, the state $x$ is asymptotically no farther than a distance $r$ from $\tilde{x}$. The quantity $r$ can be reduced with a pre-feedback controller that is applied before feedback optimization is used. More precisely, if the pair $(A,B)$ is controllable, then using pole placement we can use any $\eta>0$ to select $\rho$ such that $\rho\geq M\left\|B\right\|d_{\mathcal{U}}\eta^{-1}(2-\exp(-\rho\tau_{c,\min})+q^{\frac{\ell}{2}})$. This $\rho$ ensures $r\leq\eta$ and hence that $\|x(t,j)\|_{B_{\eta}(\tilde{x})}\to 0$ for any desired $\eta>0$.
@@ V-D Global Convergence and Robustness
Consider the hybrid system $\mathcal{H}_{FO}$ from (15) and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (7), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\big(0,\frac{2}{\lambda_{min}(Q_{u})+L}\big)$, where $Q_{u}$ is from (7) and $L:=\lambda_{max}(Q_{u}+H^{\top}Q_{y}H)$. For each maximal solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)$, for each $(t,j)\in\textnormal{dom }\phi$,
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)\|\phi(0,0)\|_{\mathcal{A}}\\ +\frac{M^{2}\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(2-\exp(-2\rho\tau_{c,\max})+q^{\frac{\ell}{2}}\Big)\exp({-\rho}t)\\ -\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(1+q^{\frac{\ell}{2}}\exp({\rho}\tau_{c,\min})\Big)\exp({-\rho}t),$ (69)
@@
which happens precisely because no computations have been performed to change the input. This case is captured by using the same steps to reach (55) with $\alpha(p)\geq\ell$ for all $p\geq 2$, using (59), and combining integrals we find
$\left\|x(t,j)-\tilde{x}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\int_{0}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(2\right)}+2}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(2\right)}+2}-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}\\ +\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(2\right)}+2}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|q^{\frac{\ell}{2}}d_{\mathcal{U}}\\ +\int_{0}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}},$ (71)
where the first integral does not have a $q^{\frac{\ell}{2}}$ term to account the zero gradient descent iterations that are performed before the first jump in $u$.
Using (57) in (71) and integrating gives
$\left\|x(t,j)-\tilde{x}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}q^{\frac{\ell}{2}}}{\rho}\big[1-\exp({-\rho}(t-t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(2\right)}+2}))\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(2\right)}+2})\big].$ (72)
Next, using $\tau_{c,\min}\leq t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(2\right)}+2}\leq 2\tau_{c,\max}$ and $\exp({-\rho}(t-\tau_{c,\min})=\exp({\rho}\tau_{c,\min})\exp({-\rho}t)$ gives the upper bound
$\left\|x(t,j)-\tilde{x}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}q^{\frac{\ell}{2}}}{\rho}\big[1-\exp({\rho}\tau_{c,\min})\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-2\rho}\tau_{c,\max})\big].$ (73)
The result follows from combining (64), (65), and (73).
@@
$C_{\iota}:=\{\zeta\in\mathcal{X}:\tau_{c}\in[0,\tau_{c,\max}+\theta_{c,\max}],\\ \tau_{g}\in[0,\tau_{g,comp}+\theta_{g,comp}]\},$ (74)
where $\theta_{c,\max}\in(-\tau_{c,\max},\infty)$ and $\theta_{g,comp}\in(-\tau_{g,comp},\infty)$. These perturbations allow $\tau_{c}$ to take values larger than $\tau_{c,\max}$ and similar for $\tau_{g}$. Although timers may be reset to inaccurate values, jumps are still triggered when at least one timer reaches zero and hence we use $D_{\iota}:=D$ in the perturbed system model.
@@
$F_{\iota}(\zeta):=\left(\begin{array}[]{cc}\big(A+\hat{A}\big)x+\big(B+\hat{B}\big)u\\ 0\\ 0\\ 0\\ -1+\kappa_{c}\\ -1+\kappa_{g}\end{array}\right),$ (75)
@@
For $G_{1,\iota}$ we have
$G_{1,\iota}(\zeta):=\left(\begin{array}[]{cc}x\\ u\\ y_{s}\\ \Pi_{\mathcal{U}}\left[z-\gamma\nabla_{u}\Phi(z,y_{s})\right]\\ \tau_{c}\\ \tau_{g,comp}+\theta_{g,comp}\end{array}\right),$ (77)
where $\theta_{g,comp}$ is from (74). This perturbed jump map allows for $\tau_{g}$ to be reset to values other than $\tau_{g,comp}$ after a gradient descent iteration is performed. For $G_{2,\iota}$ we have
$G_{2,\iota}(\zeta):=\left(\begin{array}[]{cc}x\\ z\\ \big(H+\hat{H}\big)u+d\\ z\\ \left[\tau_{c,\min}+\theta_{c,\min},\tau_{c,\max}+\theta_{c,\max}\right]\\ \tau_{g}\end{array}\right).$ (78)
Here, $\hat{H}\in\mathbb{R}^{p\times m}$ models perturbations to the matrix $H$, including those that come from the errors $\hat{A}$ and $\hat{B}$ as described above, as well as errors in the output map $\Psi$. The interval to which $\tau_{c}$ is reset is perturbed with constants $\theta_{c,\min}\in(-\tau_{c,\min},\infty)$ and $\theta_{c,\max}\in(-\tau_{c,\max},\infty)$ that satisfy $0<\tau_{c,\min}+\theta_{c,\min}\leq\tau_{c,\max}+\theta_{c,\max}$, which ensures that $\tau_{c}$ is reset to a non-empty set, though both endpoints of the interval can be perturbed. Finally we have $G_{3,\iota}(\zeta):=G_{1,\iota}(\zeta)\cup G_{2,\iota}(\zeta)$. We define
$\iota=\max\{\theta_{g,comp},\|\hat{A}x\|,\|\hat{B}u\|,\|\hat{H}u\|,\\ \kappa_{c},\kappa_{g},\theta_{c,\min},\theta_{c,\max}\!\}$ (79)
@@
$\mathcal{H}_{FO}^{\iota}=\begin{cases}\dot{\zeta}\in F_{\iota}(\zeta)&\zeta\in C_{\iota}\\ \zeta^{+}\in G_{\iota}(\zeta)&\zeta\in D_{\iota}\end{cases}.$ (80)
@@
Definition 2 ($(\tau,\epsilon)$-closeness [10, Definition 5.23]).
@@
Consider the hybrid system $\mathcal{H}_{FO}^{\iota}$ with $\iota$ as defined in (79), and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (7), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$. Then, for every $\epsilon>0$ and $\tau>0$, there exists $\delta>0$ with the following property: for every solution $\phi_{\delta}$ to $\mathcal{H}_{FO}^{\delta\iota}$, there exists a solution $\phi$ to $\mathcal{H}_{FO}$ such that $\phi_{\delta}$ and $\phi$ are $(\tau,\epsilon)$-close.
@@
Since $\mathcal{H}_{FO}$ is well-posed and its maximal solutions are complete from all initial conditions, the result follows from [24, Proposition 6.34].
For perturbations in the model and timers, Theorem 3 only holds over bounded hybrid time horizons, so the result is applied up until a chosen time $\tau$. Then for some error $\epsilon$, there is a nonzero perturbation $\delta\iota$ of the perturbed system $\mathcal{H}_{FO}^{\delta\iota}$ such that its solution $\phi_{\delta}$ is $(\tau,\epsilon)$-close to the solutions of the unperturbed system $\mathcal{H}_{FO}$.